# Revisiting the Equivalence of Bayesian Neural Networks and Gaussian Processes: On the Importance of Learning Activations

### v2

@@ 4.3 Can Learned Activations Match Performance of Closed-form Ones?
Deriving suitable neural activations to match BNNs to GPs is a formidable challenge. For example, Meronen et al., (2020) derived recently an analytical activation function for the popular Matérn kernel. However, their solution relies on the assumption that the priors on the BNN weights follow binary white noise. Such the prior hinders existing posterior learning algorithms, making the proposed solution impractical. Nevertheless, Meronen et al., (2020) shows that their solution can work with more convenient Gaussian priors, albeit as a suboptimal approximation. In this work, we propose a general gradient-based alternative that can also be applied to the GPs with the Matérn kernel and we demonstrate that our black-box approach can match or even surpass the performance of the aforementioned approximation.
Figure 3: Posterior predictive distributions for a BNN with trained parameter priors and activations (ours; 2nd row) and for a BNN with analytically derived activations (3rd and 4th row). The first column illustrates class probabilities, the second column shows the total variance in class predictions, and the last column depicts the epistemic uncertainty component of the total uncertainty. In rows 1, 2 and 4, we show posteriors obtained using HMC, while in row 3, we provide results obtained using the original code from Meronen et al., (2020) (which uses MC-Dropout instead of MCMC).
The empirical evaluation we performed for the 2D data classification problem following the setting of Meronen et al., (2020). For our method, we used a BNN with Gaussian priors with trained variances, where the activation function was modeled by an NN with a single hidden layer consisting of 5 neurons using SiLU activation. For the baseline, we trained just variances and used the fixed activation. Posterior distributions were generally obtained using a HMC sampler.
The results presented in Fig. 3 (and extended results provided in Sec. B.4 in the Appendix) demonstrate that our method enhances the match between the posteriors of a BNN and the desired target GP. Note that Meronen et al., (2020) originally used MC Dropout to obtain the posteriors, achieving much worse results than those obtained with HMC. For fairness in comparison, we tested both MC Dropout and HMC. In terms of performance, our model not only captures class probabilities accurately but also adeptly handles the total variance in class predictions and the epistemic uncertainty component, which are crucial for robust decision-making under uncertainty.
@@ 4.5 Is Model Selection for BNNs with Transferred Priors Possible?
A BNN with a functional prior learned from a GP essentially inherits the properties dictated by the GP, including those determined by specific values of the GP’s hyperparameters. While the hyperparameters of the source GP can be tuned and optimized, the BNN does not inherently possess this capability, requiring researchers to either pre-optimize the hyperparameters or look for suitable workarounds (Tran et al., , 2022). Conditioning the priors on weights, biases, and activation, as explained in Sec. 3.4, addresses this limitation, ultimately closing the BNN-GP equivalence gap. Fig. 5 shows results for a prior trained to adapt to a varying lengthscale of a Matérn kernel. The prior closely follows the behavior of the ground truth model.
@@ 4.6 Does it scale?
Learning posteriors for Gaussian Processes (GPs) is challenging, as exact inference scales cubically with the number of data points Rasmussen and Williams, (2005). To address this, scalable approximations such as Stochastic Variational Gaussian Processes (SVGP) (Hensman et al., , 2015) have been proposed for practical applications. SVGP employs stochastic variational gradients to optimize a fixed set of inducing points, typically (and also here) $1024$.
@@
Due to lack of available implementation, we failed to reproduce results from Wang et al., (2019) and provide here values copied directly from the paper.
On the other hand, there exist a number of posterior inference methods for BNNs. For the experiments presented in the previous sections, we utilized HMC. However, Variational Inference (VI) with Normalizing Flows (Rezende and Mohamed, , 2015) offers a scalable alternative, albeit with challenges related to both the accuracy of posterior approximation and selection of hyperparameters. However, the former limitation can be mitigated by employing RealNVPs (Dinh et al., , 2016) with appropriate capacity for approximating posteriors. Agrawal and Domke, (2024) showed that RealNVPs can exhibit sufficient fidelity to match posteriors for models with complex priors.
@@
where we estimate the Evidence Lower Bound (ELBO) objective for Stochastic Variational Inference (SVI) (Hoffman et al., , 2013) using minibatches $\mathcal{B}\subset\mathcal{D}$ with $|\mathcal{B}|=10{,}240$. The expectation is approximated with $S=128$ reparameterized Monte Carlo samples, where $q(w|\zeta)$ is the posterior approximation modeled by a RealNVP.
Tab. 2 summarizes the test set performance on the UCI Household Power Consumption (HouseElectric) regression dataset. This dataset contains approximately 2M rows, split into training (7/9) and test (2/9) subsets, and is the largest dataset used to evaluate GPs (Wang et al., , 2019).
@@
Variational Inference (VI) often performs better with $\beta<1$ (Higgins et al., , 2017), which also holds in our case - setting $\beta=0.4$ (a typical value) improved both RMSE and NLL. However, as indicated by discrepancies with the Exact GP results, this improvement is likely due to the reduced influence of the learned GP prior, which may be suboptimal for this particular dataset.
@@
Figure 6: Impact of the architecture of $q(w|\zeta)$ parameterized by RealNVP (Dinh et al., , 2016): (1) although adding more layers improves performance, the gains become marginal for $>16$ layers; (2) wider flows require fewer layers to achieve performance saturation.
5 Related Work
The relationship between (B)NNs and GPs has been a subject of significant research interest. Much of this work has been motivated by the desire to better understand NNs through their connection to GPs. A seminal result by Neal, (1996); Williams, (1996), later extended to deep NNs by Lee et al., (2017); Matthews et al., (2018), shows that infinitely wide layers in NNs exhibit behavior akin to GPs. While we explore a similar setting, our focus is the inverse: implementing function-space priors, specifically GP-like, within BNNs.
This problem was previously addressed by Meronen et al., (2020), who proposed an activation function for BNNs that corresponds to the popular Matérn kernel. Due to the complexity of deriving analytical solutions, other works, such as Flam-Shepherd et al., (2017, 2018); Tran et al., (2022), have explored gradient-based optimization to learn priors on weights and biases. However, their approach requires deep networks with complex distributions and additionally, presents a challenge in learning posterior. In contrast, we achieve high-fidelity functional priors in shallow BNNs by matching both priors on parameters and activations.
Finding accurate posteriors for deep and complex models such as BNNs is notoriously challenging due to their high-dimensional parameter spaces and complex likelihood surfaces. However, for single-hidden-layer wide BNNs, it has been recently shown that posterior sampling via Markov Chain Monte Carlo can be performed efficiently (Hron et al., , 2022). Moreover, research has demonstrated that the exact posterior of a wide BNN weakly converges to the posterior corresponding to the GP that matches the BNN’s prior (Hron et al., , 2020). This allows BNNs to inherit GP-like properties while preserving their advantages.
An alternative line of work touching efficient learning for BNNs in function space has explored variational inference. Sun et al., (2019) introduced functional variational BNNs (fBNNs) that maximize an Evidence Lower Bound defined over functions. However, Burt et al., (2020) highlighted issues with using the Kullback-Leibler divergence in function space, leading to ill-defined objectives. To overcome this, methods like Gaussian Wasserstein inference (Wild et al., , 2022) and Functional Wasserstein Bridge Inference (Wu et al., , 2024) leverage the Wasserstein distance to define well-behaved variational objectives. These approaches try to incorporate functional priors into variational objective for deep networks. In contrast, in our method transfer of priors to shallow BNNs remains independent and fully separated from an inference method.
Finally, other works worth mentioning include Meronen et al., (2021), who explored periodic activation functions in BNNs to connect network weight priors with translation-invariant GP priors. Pearce et al., (2020) derived BNN architectures to mirror GP kernel combinations, showcasing how BNNs can produce periodic kernels. Karaletsos and Bui, (2020) introduced a hierarchical model using GP for weights to encode correlated weight structures and input-dependent weight priors, aimed at regularizing the function space. Matsubara et al., (2021) proposed using ridgelet transforms to approximate GP function-space distributions with BNN weight-space distributions, providing a non-asymptotic analysis with finite sample-size error bounds. Finally, Tsuchida et al., (2019) extended the convergence of NN function distributions to GPs under broader conditions, including partially exchangeable priors. A more detailed discussion of related topics can be found for example, in Sections 2.3 and 4.2 of (Fortuin, , 2022).
@@ 6 CONCLUSION
Acknowledgements
@@
Appendix A Learned GP kernels
@@
Appendix B Experimental Details
B.1 Can Learned Activations Achieve More Faithful Function-space priors? -- Additional Figures
input space dimensionality = 1
input space dimensionality = 16
Figure 7: Quality of matching BNNs to the prior of a GP with a Matérn kernel ($\nu=5/2$, $\ell=1$) for 1D (left) and 16D inputs (right). We evaluate models with trained parameter priors (denoted by w), activations (denoted by a), and both (denoted by a+w). Each label specifies whether a fixed activation (e.g., ReLU) or a specific activation model (e.g., Rational) was employed. Gaussian parameter priors were used by default. If not trained, we set variances to $1$, and for the hidden layer, we normalized the variance by its width. The label w: Matérn refers to a BNN with the closed-form (fixed) activation as derived by Meronen et al., (2020).
@@ B.1 Can Learned Activations Achieve More Faithful Function-space priors? -- Additional Figures
B.2 1-dimensional Regression - Comparison with Tran et al., (2022)
In order to illustrate the abilities of our approach on a standard regression task and for a fair comparison with another method optimizing the Wasserstein distance, we followed the exact experimental setting presented in Tran et al., (2022). We used a GP with an RBF kernel (length scale $\ell$=0.6, amplitude=1.0) and a Gaussian likelihood (noise variance=0.1) as the ground truth.
For a baseline, we utilized the approach by Tran et al., (2022), consisting of a BNN with 3 layers and 50 neurons each, using TanH activation function. We considered all three prior realizations presented in that work: a simple Gaussian prior, a hierarchical prior, and a prior given by a Normalizing Flow. On the other hand, our model configuration consisted of a BNN with Gaussian priors centered at $\mathbf{0}$ and trained variances, where the activation function was modeled by a neural network with a single hidden layer with 5 neurons and SiLU activation. Posterior distributions were obtained using an HMC sampler.
We found that our approach allows for achieving the same or better results than the baseline by Tran et al., (2022), both in visual and numerical comparisons. We obtained better results in terms of distributional metrics as presented in Tab. 3 and Tab. 4 and getting better (visually) posterior distributions (see Fig. 1) without utilizing any computationally heavy prior realization (like, e.g., Normalizing Flow).
Table 3: Comparison of trained (function-space) priors in a 1D regression task: (Tran et al., , 2022) vs. our method. Whereas Tran et al., (2022) considered three different priors on parameters (including Gaussian, Hierarchical, and Normalizing Flow) for deep BNNs, our implementation consists of a single-hidden-layer BNN with standard Gaussian priors with learned variances and trained activation. The methods were compared using a set of distributional metrics against the ground truth provided by a GP. We compared prior distributions of functions over a range of $X$ covering regions with and without data (to account also for overconfidence far from data). The lower, the better.
@@
Table 4: Comparison of trained (function-space) posteriors in a 1D regression task: (Tran et al., , 2022) vs. our method. Whereas Tran et al., (2022) considered three different priors on parameters (including Gaussian, Hierarchical, and Normalizing Flow) for deep BNNs, our implementation consists of a single-hidden-layer BNN with standard Gaussian priors with learned variances and trained activation. The methods were compared on RMSE and NLL of test data.
@@
B.3 UCI regression - Comparison with Tran et al., (2022)
In addition, we compare against Tran et al., (2022) also on UCI regression tasks, which have more input dimensions $d$- between $8$ and $12$, while having 1-dimensional output. We use 10-split of training datasets as in typical in this direction of research.
For the baseline, we followed the exact experimental setting presented in Tran et al., (2022). We used a GP with an RBF kernel (length scale $\ell=\sqrt{2.0*input\_dim}$ and amplitude=1.0, $ARD$) and a Gaussian likelihood as the ground truth. We set the noise variance value and architectures for the specific datasets as in the mentioned baseline paper.
@@
Table 5: Extended results for a set of UCI regression tasks (various input dimensionality $d$) with prior transferred from a GP; comparison between the baseline (using a deep BNN; (Tran et al., , 2022)) and ours (single hidden layer BNN with 128 neurons; w/o regularization) and an activation realized by a NN with SiLU own activation were used.
@@
B.4 Can Learned Activations Match Performance of Closed-form Ones? -- Comparison with Meronen et al., (2020)
For the comparison against Meronen et al., (2020), we closely followed their experimental setting and used a GP with a Matérn kernel ($\nu=5/2$, $\ell$=1) as the ground truth. For a baseline, we utilized the approach by Meronen et al., (2020), where the authors derived analytical activations to match the Matérn kernel. Our model configuration consists of a BNN with Gaussian priors and trained variances, where the activation function was modeled by a neural network with a single hidden layer consisting of 5 neurons using SiLU activation. Posterior distributions were generally obtained using an HMC sampler, except in the case of Meronen et al., (2020), where the original implementation employed MC Dropout to approximate the posterior. However, for completeness and fairness in comparison, we also generated results using an HMC-derived posterior for the model by Meronen et al., (2020). Additional tests (see Tab. 6) were conducted on BNNs with fixed parameter priors (Default=Gaussian priors with a variance of 1, and Normal=Gaussian priors normalized by the hidden layer width) and with activation functions including ReLU and TanH.
Table 6: Similarity of the posterior predictive distributions of BNNs with a single wide hidden layer to the posterior of a GP, considered as the ground truth. Calculations were performed for a 2D test grid $X$ ($100\times 100$) spanning the area visible in Fig. 3, which includes regions both with and without training data to account for overconfidence far from the data. The posteriors were obtained using the HMC sampler in all cases, except for (Meronen et al., , 2020), where the original code was used (however, results for this model with an HMC-derived posterior are also included below). Lower values indicate better performance. For our method, we present results using weights and priors pretrained for 1D inputs, as well as those trained on functional priors for 1D/2D inputs matched to the task on $X$.
@@
B.5 Periodic and Conditional Activations

### v3

@@ 4.3 Can Learned Activations Match Performance of Closed-form Ones?
Deriving suitable neural activations to match BNNs to GPs is a formidable challenge. For example, Meronen et al. [2020] derived recently an analytical activation function for the popular Matérn kernel. However, their solution relies on the assumption that the priors on the BNN weights follow binary white noise. Such the prior hinders existing posterior learning algorithms, making the proposed solution impractical. Nevertheless, Meronen et al. [2020] shows that their solution can work with more convenient Gaussian priors, albeit as a suboptimal approximation. In this work, we propose a general gradient-based alternative that can also be applied to the GPs with the Matérn kernel and we demonstrate that our black-box approach can match or even surpass the performance of the aforementioned approximation.
Figure 5: Posterior predictive distributions for a BNN with trained parameter priors and activations (ours; 2nd row) and for a BNN with analytically derived activations (3rd row). The first column illustrates class probabilities, the second column shows the total variance in class predictions, and the last column depicts the epistemic uncertainty component of the total uncertainty. Here, we show posteriors obtained using HMC, while in Sec. D.4 of the Appendix, we provide an extended version of the figure that includes results obtained using the original code from Meronen et al. [2020] (which uses MC-Dropout instead of MCMC), along with additional numerical results.
The empirical evaluation we performed for the 2D data classification problem following the setting of Meronen et al. [2020]. For our method, we used a BNN with Gaussian priors with trained variances, where the activation function was modeled by an NN with a single hidden layer consisting of 5 neurons using SiLU activation. For the baseline, we trained just variances and used the fixed activation. Posterior distributions were generally obtained using a HMC sampler.
The results presented in Fig. 5 (and extended results provided in Sec. D.4 in the Appendix) demonstrate that our method enhances the match between the posteriors of a BNN and the desired target GP. Note that Meronen et al. [2020] originally used MC Dropout to obtain the posteriors, achieving much worse results than those obtained with HMC. For fairness in comparison, we tested both MC Dropout and HMC. In terms of performance, our model not only captures class probabilities accurately but also adeptly handles the total variance in class predictions and the epistemic uncertainty component, which are crucial for robust decision-making under uncertainty.
@@ 4.5 Is Model Selection for BNNs with Transferred Priors Possible?
A BNN with a functional prior learned from a GP essentially inherits the properties dictated by the GP, including those determined by specific values of the GP’s hyperparameters. While the hyperparameters of the source GP can be tuned and optimized, the BNN does not inherently possess this capability, requiring researchers to either pre-optimize the hyperparameters or look for suitable workarounds [Tran et al., 2022]. Conditioning the priors on weights, biases, and activation, as explained in Sec. 3.4, addresses this limitation, ultimately closing the BNN-GP equivalence gap. Fig. 7 shows results for a prior trained to adapt to a varying lengthscale of a Matérn kernel. The prior closely follows the behavior of the ground truth model.
@@ 4.6 Does It Scale?
Learning posteriors for Gaussian Processes (GPs) is challenging, as exact inference scales cubically with the number of data points Rasmussen and Williams [2005]. To address this, scalable approximations such as Stochastic Variational Gaussian Processes (SVGP) [Hensman et al., 2015] have been proposed for practical applications. SVGP employs stochastic variational gradients to optimize a fixed set of inducing points, typically (and also here) $1024$.
@@
Due to lack of available implementation, we failed to reproduce results from Wang et al. [2019] and provide here values copied directly from the paper.
On the other hand, there exist a number of posterior inference methods for BNNs. For the experiments presented in the previous sections, we utilized HMC. However, Variational Inference (VI) with Normalizing Flows [Rezende and Mohamed, 2015] offers a scalable alternative, albeit with challenges related to both the accuracy of posterior approximation and selection of hyperparameters. However, the former limitation can be mitigated by employing RealNVPs [Dinh et al., 2016] with appropriate capacity for approximating posteriors. Agrawal and Domke [2024] showed that RealNVPs can exhibit sufficient fidelity to match posteriors for models with complex priors.
@@
where we estimate the Evidence Lower Bound (ELBO) objective for Stochastic Variational Inference (SVI) [Hoffman et al., 2013] using minibatches $\mathcal{B}\subset\mathcal{D}$ with $|\mathcal{B}|=10{,}240$. The expectation is approximated with $S=128$ reparameterized Monte Carlo samples, where $q(w|\zeta)$ is the posterior approximation modeled by a RealNVP.
Tab. 2 summarizes the test set performance on the UCI Household Power Consumption (HouseElectric) regression dataset. This dataset contains approximately 2M rows, split into training (7/9) and test (2/9) subsets, and is the largest dataset used to evaluate GPs [Wang et al., 2019].
@@
Variational Inference (VI) often performs better with $\beta<1$ [Higgins et al., 2017], which also holds in our case - setting $\beta=0.4$ (a typical value) improved both RMSE and NLL. However, as indicated by discrepancies with the Exact GP results, this improvement is likely due to the reduced influence of the learned GP prior, which may be suboptimal for this particular dataset.
@@
Figure 8: Impact of the architecture of $q(w|\zeta)$ parameterized by RealNVP [Dinh et al., 2016]: (1) although adding more layers improves performance, the gains become marginal for $>16$ layers; (2) wider flows require fewer layers to achieve performance saturation.
@@ 5 Conclusion
{contributions}
Marcin Sendera actively participated in shaping the project across all its stages, contributing both during discussions and through hands-on development. He explored and assessed multiple variants of the Wasserstein loss, as well as a range of learnable activation functions proposed in the literature, including piecewise linear (PWL), Padé, and others. He proposed to use the closed-form solution for the Wasserstein metric. He was primarily responsible for implementing and executing the experiments reported in Figures 3 and 4, and in Table 1, ensuring the reproducibility and robustness of the results. He contributed to writing of the manuscript, particularly in sections related to empirical results.
Amin Sorkhei participated in discussions at the initial stage of the project. He implemented and conducted the initial experiments involving the mapping of Gaussian Process priors to Bayesian Neural Networks using the Kullback-Leibler divergence and a learnable activation parameterized by a neural network.
Tomasz Kuśmierczyk conceived the project and supervised its development across all stages, including the core theoretical contributions and experimental design. He introduced the idea of learning activations and proposed the use of Wasserstein-based loss and regularizations. He proposed to use neural networks to parameterize learnable activations. He developed and implemented the periodic (Section 3.3) and conditioned activations (Section 3.4). He was responsible for implementing and executing the experiments presented in Figures 2, 5, 6, 7, 8, and Table 2. He was responsible for the majority of the manuscript writing as well as for shaping its final form.
Acknowledgements.
We thank Nikolay Malkin and Jacek Tabor for helpful discussions at early stages of this project and revising the manuscript.
@@ Acknowledgements.
Appendix A Related Work
The relationship between (B)NNs and GPs has been a subject of significant research interest. Much of this work has been motivated by the desire to better understand NNs through their connection to GPs. A seminal result by Neal [1996], Williams [1996], later extended to deep NNs by Lee et al. [2017], Matthews et al. [2018], shows that infinitely wide layers in NNs exhibit behavior akin to GPs. While we explore a similar setting, our focus is the inverse: implementing function-space priors, specifically GP-like, within BNNs.
This problem was previously addressed by Meronen et al. [2020], who proposed an activation function for BNNs that corresponds to the popular Matérn kernel. Due to the complexity of deriving analytical solutions, other works, such as Flam-Shepherd et al. [2017, 2018], Tran et al. [2022], have explored gradient-based optimization to learn priors on weights and biases. However, their approach requires deep networks with complex distributions and additionally, presents a challenge in learning posterior. In contrast, we achieve high-fidelity functional priors in shallow BNNs by matching both priors on parameters and activations.
Finding accurate posteriors for deep and complex models such as BNNs is notoriously challenging due to their high-dimensional parameter spaces and complex likelihood surfaces. However, for single-hidden-layer wide BNNs, it has been recently shown that posterior sampling via Markov Chain Monte Carlo can be performed efficiently [Hron et al., 2022]. Moreover, research has demonstrated that the exact posterior of a wide BNN weakly converges to the posterior corresponding to the GP that matches the BNN’s prior [Hron et al., 2020]. This allows BNNs to inherit GP-like properties while preserving their advantages.
An alternative line of work touching efficient learning for BNNs in function space has explored variational inference. Sun et al. [2019] introduced functional variational BNNs (fBNNs) that maximize an Evidence Lower Bound defined over functions. However, Burt et al. [2020] highlighted issues with using the Kullback-Leibler divergence in function space, leading to ill-defined objectives. To overcome this, methods like Gaussian Wasserstein inference [Wild et al., 2022] and Functional Wasserstein Bridge Inference [Wu et al., 2024] leverage the Wasserstein distance to define well-behaved variational objectives. These approaches try to incorporate functional priors into variational objective for deep networks. In contrast, in our method transfer of priors to shallow BNNs remains independent and fully separated from an inference method.
Finally, other works worth mentioning include Meronen et al. [2021], who explored periodic activation functions in BNNs to connect network weight priors with translation-invariant GP priors. Pearce et al. [2020] derived BNN architectures to mirror GP kernel combinations, showcasing how BNNs can produce periodic kernels. Karaletsos and Bui [2020] introduced a hierarchical model using GP for weights to encode correlated weight structures and input-dependent weight priors, aimed at regularizing the function space. Matsubara et al. [2021] proposed using ridgelet transforms to approximate GP function-space distributions with BNN weight-space distributions, providing a non-asymptotic analysis with finite sample-size error bounds. Finally, Tsuchida et al. [2019] extended the convergence of NN function distributions to GPs under broader conditions, including partially exchangeable priors. A more detailed discussion of related topics can be found for example, in Sections 2.3 and 4.2 of [Fortuin, 2022].
Appendix B Computational and Architectural Advantages
Our method involves an initial prior fitting step, which precedes the posterior inference step. The prior fitting step takes approximately one hour for the HouseElectric dataset (Section 4.6); however, one can alternatively use a prior from a library of pre-trained priors, thereby amortizing the cost of this step.
Our experiments show that substituting fixed activations (ReLU/TanH, $\sim 0.5$s/iteration) with learnable NN activations (single hidden layer, 5 neurons, $\sim 2.3$s/iteration) or periodic activations ($\sim 2.7$s/iteration) increases the time per iteration by approximately a factor of 4 on our 16-core CPU setup. On the other hand, learning parameter priors alongside activations has only a marginal effect on computation time compared to fixing priors and only learning activations.
Additionally, we conducted a study measuring both the quality of prior fit and the runtime for a prior modeled by a neural network with varying numbers of layers and widths. The respective results can be found in Figures 1010 and Tables 34. These results suggest that increasing the number of layers up to a certain depth (e.g., 4 layers in this case) may marginally improve the fidelity of the solutions, whereas increasing the network width tends to make convergence harder. In either case, using a larger neural network consistently leads to longer convergence times.
Figure 9: Prior matching quality: fixed width.
Figure 10: Prior matching quality: fixed depth.
Table 3: Prior matching time
(activation=NN, width=5)
#layers iteration time [s]
1 2.34
2 3.10
4 4.57
8 7.38
16 15.03
Table 4: Prior matching time
(activation=NN, single hidden layer)
#neurons iteration time [s]
5 2.34
10 2.59
20 3.95
40 7.59
Once the prior is established (or a pre-trained one is used), posterior inference can be performed. Our approach offers here computational advantages, particularly for large datasets where exact GP inference is prohibitive. For example, RealNVP-based variational inference (Section 4.6) converges in about 2 hours on a single GPU for HouseElectric, while exact GP inference can take over three days on a single GPU, or 1.5 hours when parallelized on 8 GPUs using methods such as BBMM [Wang et al., 2019]. Furthermore, during posterior inference using HMC for the data in Fig. 3, we observed no significant differences in step time ($\sim 1$s) between fixed and learnable activations, though this may vary with implementation and hardware.
Finally, beyond computational performance, BNNs provide greater modularity and extensibility. Single-layer BNNs can serve as components in larger neural architectures, for example, as Bayesian last layers (with remaining layers trained point-wise). This enables leveraging architectural innovations from deep learning in ways that GPs cannot easily accommodate. Our method could replace existing approaches in models like SNGP [Liu et al., 2020], potentially offering improved prior control (as random Fourier feature GPs are limited to specific stationary kernels), more off-the-shelf end-to-end training options (e.g., SVGD), and potentially enhanced performance, as suggested by our comparisons against SVGP in Section 4.6.
Appendix C Learned GP Kernels
@@
Appendix D Experimental Details
D.1 Can Learned Activations Achieve More Faithful Function-space Priors? – Additional Figures
@@
D.2 1-dimensional Regression - Comparison with Tran et al. [2022]
In order to illustrate the abilities of our approach on a standard regression task and for a fair comparison with another method optimizing the Wasserstein distance, we followed the exact experimental setting presented in Tran et al. [2022]. We used a GP with an RBF kernel (length scale $\ell$=0.6, amplitude=1.0) and a Gaussian likelihood (noise variance=0.1) as the ground truth.
For a baseline, we utilized the approach by Tran et al. [2022], consisting of a BNN with 3 layers and 50 neurons each, using TanH activation function. We considered all three prior realizations presented in that work: a simple Gaussian prior, a hierarchical prior, and a prior given by a Normalizing Flow. On the other hand, our model configuration consisted of a BNN with Gaussian priors centered at $\mathbf{0}$ and trained variances, where the activation function was modeled by a neural network with a single hidden layer with 5 neurons and SiLU activation. Posterior distributions were obtained using an HMC sampler.
We found that our approach allows for achieving the same or better results than the baseline by Tran et al. [2022], both in visual and numerical comparisons. We obtained better results in terms of distributional metrics as presented in Tab. 5 and Tab. 6 and getting better (visually) posterior distributions (see Fig. 3) without utilizing any computationally heavy prior realization (like, e.g., Normalizing Flow).
Table 5: Comparison of trained (function-space) priors in a 1D regression task: [Tran et al., 2022] vs. our method. Whereas Tran et al. [2022] considered three different priors on parameters (including Gaussian, Hierarchical, and Normalizing Flow) for deep BNNs, our implementation consists of a single-hidden-layer BNN with standard Gaussian priors with learned variances and trained activation. The methods were compared using a set of distributional metrics against the ground truth provided by a GP. We compared prior distributions of functions over a range of $X$ covering regions with and without data (to account also for overconfidence far from data). The lower, the better.
@@
Table 6: Comparison of trained (function-space) posteriors in a 1D regression task: [Tran et al., 2022] vs. our method. Whereas Tran et al. [2022] considered three different priors on parameters (including Gaussian, Hierarchical, and Normalizing Flow) for deep BNNs, our implementation consists of a single-hidden-layer BNN with standard Gaussian priors with learned variances and trained activation. The methods were compared on RMSE and NLL of test data.
@@
D.3 UCI Regression - Comparison with Tran et al. [2022]
In addition, we compare against Tran et al. [2022] also on UCI regression tasks, which have more input dimensions $d$ – between $8$ and $12$, while having 1-dimensional output. We use 10-split of training datasets as in typical in this direction of research.
For the baseline, we followed the exact experimental setting presented in Tran et al. [2022]. We used a GP with an RBF kernel (length scale $\ell=\sqrt{2.0*input\_dim}$ and amplitude=1.0, $ARD$) and a Gaussian likelihood as the ground truth. We set the noise variance value and architectures for the specific datasets as in the mentioned baseline paper.
@@
Table 7: Extended results for a set of UCI regression tasks (various input dimensionality $d$) with prior transferred from a GP; comparison between the baseline (using a deep BNN; [Tran et al., 2022]) and ours (single hidden layer BNN with 128 neurons; w/o regularization) and an activation realized by a NN with SiLU own activation were used.
@@
D.4 Can Learned Activations Match Performance of Closed-form Ones? – Comparison with Meronen et al. [2020]
For the comparison against Meronen et al. [2020], we closely followed their experimental setting and used a GP with a Matérn kernel ($\nu=5/2$, $\ell$=1) as the ground truth. For a baseline, we utilized the approach by Meronen et al. [2020], where the authors derived analytical activations to match the Matérn kernel. Our model configuration consists of a BNN with Gaussian priors and trained variances, where the activation function was modeled by a neural network with a single hidden layer consisting of 5 neurons using SiLU activation. Posterior distributions were generally obtained using an HMC sampler, except in the case of Meronen et al. [2020], where the original implementation employed MC Dropout to approximate the posterior. However, for completeness and fairness in comparison, we also generated results using an HMC-derived posterior for the model by Meronen et al. [2020]. Additional tests (see Tab. 8) were conducted on BNNs with fixed parameter priors (Default=Gaussian priors with a variance of 1, and Normal=Gaussian priors normalized by the hidden layer width) and with activation functions including ReLU and TanH.
Figure 12: Posterior predictive distributions for a BNN with trained parameters priors and activations (ours; 2nd row), and for a BNN with analytically derived activation (3rd and 4th row). The first column illustrates class probabilities, the second column shows the total variance in class predictions, and the last column depicts the epistemic uncertainty component of the total uncertainty.
Table 8: Similarity of the posterior predictive distributions of BNNs with a single wide hidden layer to the posterior of a GP, considered as the ground truth. Calculations were performed for a 2D test grid $X$ ($100\times 100$) spanning the area visible in Fig. 12, which includes regions both with and without training data to account for overconfidence far from the data. The posteriors were obtained using the HMC sampler in all cases, except for [Meronen et al., 2020], where the original code was used (however, results for this model with an HMC-derived posterior are also included below). Lower values indicate better performance. For our method, we present results using weights and priors pretrained for 1D inputs, as well as those trained on functional priors for 1D/2D inputs matched to the task on $X$.
@@
D.5 Periodic and Conditional Activations